OpenStudy (anonymous):

Which of the following inequalities does the graph represent? t ≤ –3 or t ≥ 4 t < –3 or t > 4 –3 < t < 4 –3 ≤ t ≤ 4

OpenStudy (anonymous):

A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1 A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1 A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1

OpenStudy (anonymous):

graph attached

OpenStudy (anonymous):

3<=x<=4

OpenStudy (anonymous):

is it c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I meant negative three just so you know

OpenStudy (anonymous):

your welcome